# RuleR: Improving LLM Controllability by Rule-based Data Recycling

### v3

@@ Abstract
Despite the remarkable advancement of Large language models (LLMs), they still lack delicate controllability under sophisticated constraints, which is critical to enhancing their response quality and the user experience. While conditional supervised fine-tuning (SFT) can potentially improve LLM controllability, curating new SFT data to fulfill the constraints usually relies on human experts or proprietary LLMs, which is time-consuming and expensive. To bridge this gap, we propose Rule-based Data Recycling (RuleR), a human/LLM-free data augmentation method incorporating multiple constraints into the original SFT data. Instead of creating new responses from scratch, RuleR integrates linguistic or formatting rules into the original instructions and modifies the responses to fulfill the rule-defined constraints. Training on the “recycled” data consolidates LLM capability to generate constrained outputs. Extensive experiments demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following performance. RuleR’s code is released on https://github.com/tianyi-lab/RuleR.
@@
Ming Li*1, Han Chenthanks: Equal Contribution., Chenguang Wang*2, Dang Nguyen1, Dianqi Li, Tianyi Zhou1 1University of Maryland 2Stony Brook University {minglii, tianyi}@umd.edu
@@ 1 Introduction
Figure 1: Comparing widely-used data generation strategy (top) and RuleR (bottom) enhancing LLM controllability. Most existing methods rely on human/model supervision or reflection to generate new instructions and responses. However, discarding existing data is a waste of efforts taken to curate them by human experts or expert LLMs. On the contrary, our RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
Figure 2: Examples of the data recycling workflows in Rule-based Data Recycling (RuleR). (a), (b) and (c) select different predefined rules to modify the original data to fulfill constraints on the complexity or format of the response. The differences in new responses are highlighted in red, the example in (a) has already satisfied the appended constraint, thus the response is kept unchanged.
@@
Despite the remarkable advancement of the current Large language models (LLMs) and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b); Jiang et al. (2023), one critical challenge is to generate responses better interacting with humans, with the utility and effectiveness maximized for end-users Liu et al. (2024). According to the systematic investigation from Liu et al. (2024), it is essential for LLMs to constrain their outputs to follow user-specified formats or characteristics. In various practical applications, without any constraint or format being enforced, free-formed responses are not legal or directly applicable. It has also been verified on LLM Agents Li et al. (2024f); Chen et al. (2023, 2024b); Zhang et al. (2024) that enforcing predefined formats is necessary for various tasks.
@@
However, existing SFT datasets are mainly composed of general instructions without user-specified constraints Wei et al. (2022); Wang et al. (2022); Taori et al. (2023); Xu et al. (2023); Zhou et al. (2023a); Li et al. (2023a); Zhang et al. (2023); Xu et al. (2024) and thus result in models lacking delicate controllability of the lengths and format of responses Chen et al. (2024a); Xia et al. (2024). To enhance the utility of existing SFT data in improving the controllability of LLMs, a potential method is to rewrite or modify instructions and responses by experts such as humans/LLMs Xu et al. (2023); Li et al. (2023a, 2024b, 2024a); He et al. (2024) in order to make them fulfilling multiple constraints, as shown in Figure 1 (top). However, the curation of new data is not only costly and inefficient, requiring careful editing by human experts or proprietary LLMs, but also represents a waste of previous efforts: It is impractical to discard all existing data and create brand new data every time we need to add more constraints to the instructions. Hence, we raise the question: Can we “recycle” existing SFT data without human/LLM editing and enforce various types of constraints in order to improve LLM controllability?
@@
Drawing inspiration from IFEval Zhou et al. (2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, and the human/model-free data augmentation in Mosaic-IT Li et al. (2024c), we propose Rule-based Data Recycling (RuleR), which automatically “re-engineering” existing SFT data for improving LLM controllability. As illustrated in Figure 1 (bottom), the key insight of RuleR is to automatically build constraint-augmented SFT dataset at no cost of human/LLM, by applying predefined rules to enforce constraints in the instructions and responses of the original dataset. Specifically, we manually inspect and construct a diverse set of rules as constraints, which specify the linguistic or formatting constraints on different parts of the response. To modify the original dataset, these rules are appended to the instructions, while the corresponding responses are modified if necessary to fulfill the constraints of these rules.
@@
Our predefined rules cover a wide range of diverse constraints generalizable to many application scenarios, ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats of some keywords. Each rule is composed of (1) multiple templates to produce additional instructions enforcing the constraints, and (2) a piece of code [^1] https://github.com/tianyi-lab/RuleR that alternately edits the original instruction and response in order to make the edited response fulfill all the constraints appended to the instruction. For each sample from the original dataset, we randomly draw several rules to be applied to the editing. This produces an augmented sample with the constraints enforced so it can be used for controllability tuning. The complete list of rules and descriptions can be found in Appendix C and Table 4.
@@
Illustrative examples are provided in Figure 2, which showcase (a) a rule that constrains the number of letters of the response; (b) a rule that specifies the case of specific words (if presenting in the response); and (c) a rule that specifies the wrapping format of specific words. To ensure the consistency between the input constraints and the output response, we modify both the instruction and response based on the characteristics of the original data sample. For each sample, we only sample from the rules applicable to the original response, hence avoiding the potential incorrectness of the edited responses.
@@
Extensive experimental results on the IFEval (Instruction-Following Eval) Zhou et al. (2023b) benchmark on various base models and datasets demonstrate the effectiveness of RuleR in enhancing LLM controllability without extra help from human/model. On the other hand, RuleR still preserves the general instruction-following ability promoted by the original SFT dataset. This is demonstrated by the instruction-following metrics (Pair-wise comparison and Open LLM Leaderboard). To the best of our knowledge, RuleR is the first human/model-free data augmentation and recycling approach designed to improve LLM controllability under multiple constraints.
@@ 2.1 Preliminaries
[Instruction]
Ensure the 2nd paragraph is enclosed in "-| |-" in the response.
[Response]
[Instruction]
Keep the word "way" bounded by "@ @" in the response if it appears.
Describe an example of a time you used to influence in a positive way
[Response]
@@ 2.2.1 Rule Construction
While most existing methods still require human experts or strong teacher LLMs to generate new data (Figure 1 (top)), we aim at “re-engineering” instructions and responses from existing SFT datasets for building controllability-focused datasets, without expensive and time-consuming supervision from humans or LLMs. In the following, we introduce a rule-based approach “Rule-based Data Recycling (RuleR)” to create high-quality data augmentations for improving LLM controllability.
RuleR reformulates the original instructions and responses by applying rule-based edits according to pre-defined constraints. However, not every constraint can be applied to each sample without humans or LLMs. Hence, we only incorporate constraints compatible with the original sample and those can be implemented with simple rectifications like regular expressions. In RuleR, the rules and constraints are elaborately designed and selected to specify or change certain characteristics of the original data, resulting in consistency between each constraint-augmented instruction and the reformulated response.
Specifically, we focus on the characteristics of responses that can be defined by rules, e.g., by checking the 220 distinct linguistic features in the LFTK package Lee and Lee (2023). In addition, we collect constraints from existing works and widely used instruction-tuning datasets. These diverse characteristics include punctuation-, word-, sentence-, paragraph-level occurrences, and frequencies. We construct rules constraining or specifying the characteristics of the original response. In conjunction with the rules containing these characteristics, we also create rules specifying the format of responses to improve LLM’s format-following capability.
To ensure that the rules selected or sampled for each sample are applicable and legal, we apply the following additional protocols: (1) The rules need to be applicable to the original sample. For instance, the rule “Generating a title before giving the response” is not applicable as we can not generate a title without the help of humans or other additional models. (2) The rules should not remove the content of the original response. For instance, the rule “Ensure the word xxx is not shown in the response” is not applicable since we can not directly remove this word from the response as it is part of the original response. (3) The rules should be compatible with the original sample. If the rule is “Ensure there are more than N sentences in the response”, then it can not be applied to samples whose responses have $<N$ sentences. The complete list of rules is provided in Table 4, which covers both high-level constraints such as the term frequency, and lower-level constraints such as specific wrapping formats.
@@ 2.2.2 Rule Implementation
To implement each rule as a program editing instructions and responses, we formulate each pre-defined rule as a tuple, $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated templates for creating instructions of the $k^{th}$ rule, while $f_{k}$ and $g_{k}$ are the corresponding functions to reformulate instructions and responses (if necessary), respectively.
Specifically, the function $f_{k}$ selects an appropriate template of a rule for a given data sample and augments the original instruction with an instruction generated by the template. It first selects a subset of rules applicable to the characteristics of the response (e.g., presence of keywords, number of sentences, etc.). Then, we randomly draw one rule out of the subset and create a formatted instruction of the rule from the template. Such rule instruction is appended to the original instruction.
@@
The function $g_{k}$ is designed to modify the response to be consistent with the augmented instruction (after applying function $f_{k}$), i.e., with the rule applied. Applying $g_{k}$ either preserves the original response or revises it. Some rules do not require editing of responses, e.g., keyword appearance, the number of nouns, etc. For rules defining the case or format of certain parts in the response, modifications are needed. The detailed descriptions of each predefined rule can be found in Appendix C. Specifically, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
@@ 2.2.3 Multi-rule Implementation
§ 2.2.2 illustrates how RuleR applies one rule to edit each data sample. To create more complex, diverse, and challenging samples, we can extend it to multiple rules randomly drawn from the feasible set of rules. We provide examples of multi-rule augmentation in Figure 3. It shows that the complexity of the augmented data grows dramatically, forcing the LLMs to learn to follow multiple constraints.
@@
3 Experimental Setup
3.1 Implementation Details
The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. For training on Llama2-7B and Llama2-13B Touvron et al. (2023b), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs. For training on Mistral-7B Jiang et al. (2023), the maximum learning rate is set to $1\times 10^{-5}$ with a warmup rate of $0.1$ for $2$ epochs. When utilizing our method, we run the augmentation process $3$/$2$ times to simulate the epochs of training. These augmented data are then mixed together and used for training $1$ epoch. All other configurations are kept the same as the baselines.
3.2 Training Datasets
We utilize 5 SFT datasets to evaluate the effectiveness of our method: Alpaca dataset Taori et al. (2023), Alpaca-GPT4 dataset Peng et al. (2023), WizardLM dataset Xu et al. (2023), Recycled WizardLM Dataset Li et al. (2023a), DEITA dataset Liu et al. (2023). A detailed illustration of datasets can be found in the Appendix A.
3.3 Evaluation Metrics
Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts. This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al. (2024); Li et al. (2023b); Sottana et al. (2023). We utilize test instruction sets from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023b), comprising 218 and 80 diverse human-curated instructions, respectively. Following the framework by Li et al. (2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy. To mitigate positional bias Ko et al. (2020); Wang et al. (2023a), responses are presented in two distinct sequences. Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction. One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions. The specific evaluation prompt is presented in the Appendix B.
@@ 4.1 Main Results
The main experimental results are presented in Table 1, containing the performance comparison on the Instruction Following Eval Zhou et al. (2023b), Pair-wise Comparison Winning Score, and the Open LLM Leaderboard Gao et al. (2021a), on 3 different base models and several different instruction tuning datasets. The Pair-wise Comparison Winning Score is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$ and the values that are greater than $1.0$ represent better responses generated. Detailed descriptions of evaluation metrics can be found in the Appendix B.
@@ 4.2 Ablation Studies
Table 2: Ablation Study. Average scores on “IF Eval” and “Open LLM” of Mistral-7B finetuned with ours-augmented Alpaca-GPT4 data. “(*)” represents default configuration.

### v4

@@ Abstract
Despite the remarkable advancement of Large language models (LLMs), they still lack delicate controllability under sophisticated constraints, which is critical to enhancing their response quality and the user experience. While supervised fine-tuning (SFT) can potentially improve LLM controllability, curating new SFT data to fulfill the constraints usually relies on human experts or proprietary LLMs, which is time-consuming and expensive. To bridge this gap, we propose Rule-based Data Recycling (RuleR), a human/LLM-free data augmentation method incorporating multiple constraints into the original SFT data. Instead of creating new responses from scratch, RuleR integrates linguistic or formatting rules into the original instructions and modifies the responses to fulfill the rule-defined constraints. Training on the “recycled” data consolidates LLM capability to generate constrained outputs, improving LLM controllability while maintaining promising general instruction-following capabilities.
@@
Ming Li*1, Han Chenthanks: Equal Contribution., Chenguang Wang*2, Dang Nguyen1, Dianqi Li, Tianyi Zhou1 1University of Maryland 2Stony Brook University {minglii, tianyi}@umd.edu Project: https://github.com/tianyi-lab/RuleR
@@ 1 Introduction
Figure 1: Comparing widely-used data generation strategy (top) and RuleR (bottom) enhancing LLM controllability. Most existing methods rely on human/model rewriting to generate new instructions and responses. However, discarding existing data is a waste of effort. Our RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can generate new SFT data that improves LLM controllability.
@@
Despite the remarkable advancement of the current Large language models (LLMs) and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets, one critical challenge is to generate responses better interacting with humans, with the utility and effectiveness maximized for end-users Liu et al. (2024); Huang et al. (2024, 2023). According to the systematic investigation from Liu et al. (2024), it is essential for LLMs to constrain their outputs to follow user-specified formats or characteristics. In various practical applications, free-formed responses are not legal or directly applicable without any constraint or format being enforced. It has also been verified on LLM Agents Li et al. (2024f); Chen et al. (2023, 2024b); Zhang et al. (2024) that enforcing predefined formats is necessary for tasks.
@@
However, existing SFT datasets are mainly composed of general instructions without user-specified constraints Wei et al. (2022); Wang et al. (2022); Taori et al. (2023); Xu et al. (2023); Zhou et al. (2023a); Li et al. (2023a); Zhang et al. (2023); Xu et al. (2024) and thus result in models lacking delicate controllability of the lengths and format of responses Chen et al. (2024a); Xia et al. (2024). To enhance the utility of existing SFT data in improving the controllability of LLMs, a potential method is to rewrite or modify instructions and responses by experts such as humans/LLMs Xu et al. (2023); Li et al. (2023a, 2024b, 2024a); He et al. (2024); Dong et al. (2024); Wu et al. (2024) in order to make them fulfilling multiple constraints, as shown in Figure 1 (top). However, the curation of new data is not only costly and inefficient, requiring careful editing by human experts or proprietary LLMs, but also represents a waste of previous efforts: It is impractical to discard all existing data and create brand new data every time we need to add more constraints to the instructions. Hence, we raise the question: Can we “recycle” existing SFT data without human/LLM editing and enforce various types of constraints in order to improve LLM controllability?
@@
Drawing inspiration from IFEval Zhou et al. (2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, and the human/model-free data augmentation in Mosaic-IT Li et al. (2024c), we propose Rule-based Data Recycling (RuleR), which automatically “recycles” existing SFT data for improving LLM controllability. As illustrated in Figure 1 (bottom), the key insight of RuleR is to automatically build constraint-augmented SFT datasets at no cost of human/LLM efforts, by applying predefined rules to the original instructions and responses. Specifically, we manually inspect and construct a diverse set of rules as constraints, which specify the linguistic or formatting constraints on different parts of the response.
@@
Our predefined rules cover a wide range of diverse constraints generalizable to many application scenarios, ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats of some keywords. Each rule is composed of (1) multiple templates to produce additional instructions enforcing the constraints, and (2) a piece of code that alternately edits the original instruction and response in order to make the edited response fulfill all the constraints appended to the instruction. For each sample from the original dataset, we randomly draw several rules to be applied to the editing. This produces an augmented sample with the constraints enforced so it can be used for controllability tuning. The complete list of rules and descriptions can be found in Appendix E.
@@
Illustrative examples are provided in Figure 2, which showcase (a) a rule that constrains the number of letters; (b) a rule that specifies the case of specific words (if presenting in the response); and (c) a rule that specifies the wrapping format of specific words. To ensure the consistency between the input constraints and the output response, we modify both the instruction and response based on the characteristics of the original data sample. For each sample, we only sample from the rules applicable to the original response, hence avoiding the potential incorrectness of the edited responses.
Figure 2: Examples of our data recycling workflows. (a), (b) and (c) select different predefined rules to modify the original data to fulfill constraints on the complexity or format of the response. The differences in new responses are highlighted in red, the example in (a) has already satisfied the appended constraint, thus the response is kept unchanged.
@@
Extensive experimental results on the IFEval benchmark on various base models and datasets demonstrate the effectiveness of RuleR in enhancing LLM controllability without extra help from humans/models. On the other hand, RuleR still preserves the general instruction-following ability promoted by the original SFT dataset. This is demonstrated by the instruction-following metrics (Pair-wise comparison and Open LLM Leaderboard). To the best of our knowledge, RuleR is the first human/model-free data augmentation and recycling approach designed to improve LLM controllability under multiple constraints.
@@ 2.2.1 Rule Construction
While most existing methods still require human experts or strong teacher LLMs to generate new data (Figure 1 (top)), we aim at “recycling” instructions and responses from existing SFT datasets to build controllability-focused datasets, without expensive and time-consuming supervision from humans or LLMs. In the following, we introduce a rule-based approach “Rule-based Data Recycling (RuleR)” to create high-quality augmented data for improving LLM controllability.
RuleR reformulates the original instructions and responses by applying rule-based edits according to pre-defined constraints. However, not every constraint is applicable to a randomly selected sample without fully rewriting. Hence, we only incorporate constraints compatible with the original sample and those can be implemented with simple rectifications like regular expressions. Specifically, we focus on the characteristics of responses that can be defined by rules, e.g., by checking the 220 distinct linguistic features in the LFTK package Lee and Lee (2023). In addition, we collect constraints from existing works and widely used instruction-tuning datasets. These diverse characteristics include punctuation-, word-, sentence-, paragraph-level occurrences, and frequencies. Thus we construct rules constraining or specifying the characteristics of the original response. In conjunction with the rules containing these characteristics, we also create rules specifying the format of responses to improve LLM’s format-following capability.
To ensure that the rules selected or sampled for each sample are applicable, we apply the following additional protocols: (1) The rules need to be applicable to the original sample. For instance, the rule “Generating a title before giving the response” is not applicable as we can not generate a title without the help of humans or other additional models. (2) The rules should not include removing the content of the original response. For instance, the rule “Ensure the word xxx is not shown in the response” is not applicable since we can not directly remove this word from the response as the removal might disrupt the original semantic integrity. (3) The rules should be compatible with the original sample. If the rule is “Ensure there are more than N sentences in the response”, then it can not be applied to samples whose responses have $<N$ sentences. The complete list of rules is provided in Table 4, which covers both high-level constraints such as the term frequency, and lower-level constraints such as specific wrapping formats.
@@ 2.2.2 Rule Implementation
To implement each rule to original instructions and responses, we notate each pre-defined rule as a tuple for simplicity, $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated instruction templates for creating instructions of the $k^{th}$ rule, while $f_{k}$ and $g_{k}$ are the corresponding functions to reformulate instructions and responses (if necessary), respectively.
Specifically, the function $f_{k}$ selects an appropriate template of a rule for a given data sample and augments the original instruction with an instruction generated by the template. It first selects a subset of rules applicable to the characteristics of the response (e.g., presence of keywords, number of sentences, etc.). Then, it randomly draw one rule out of the subset and create a formatted instruction of the rule from the template. Such rule instruction is appended to the original instruction.
@@
The function $g_{k}$ is designed to modify the response to be consistent with the augmented instruction (after applying function $f_{k}$), i.e., with the rule applied. Applying $g_{k}$ either preserves the original response or revises it. Some rules do not require editing of responses, e.g., keyword appearance, the number of nouns, etc. For rules defining the case or format of certain parts in the response, modifications are needed. The detailed descriptions of each predefined rule can be found in Appendix E. Specifically, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
@@ 2.2.3 Multi-rule Implementation
To create more complex, diverse, and challenging samples, we can extend the previous process to multiple rules randomly drawn from the feasible set of rules. We provide examples of multi-rule augmentation in Appendix B, which forces LLMs to learn to follow multiple constraints.
The detailed experimental setup can be found in Appendix A, including Implementation Details, Training Datasets, and Evaluation Metrics.
@@ 3.1 Main Results
The main experimental results are presented in Table 1, containing the performance comparison on the Instruction Following Eval Zhou et al. (2023b), Pair-wise Comparison Winning Score, and the Open LLM Leaderboard Gao et al. (2021a), on 3 different base models and several different instruction tuning datasets. The Pair-wise Comparison Winning Score is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$ and the values that are greater than $1.0$ represent better responses generated. Detailed descriptions of evaluation metrics can be found in the Appendix D.
@@ 3.2 Ablation Studies
Table 2: Ablation Study. “(*)” represents default.
In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset, aiming to evaluate the impact of several factors.
Effect of Templates: As shown in Table 2, “Single Temp.” represents utilizing only one rule-instruction template for each rule, while “Diverse Temp.” represents utilizing several different templates, approximately $10$, with the same meaning for each rule and those templates are randomly sampled during the augmentation. “Diverse Temp.” demonstrates superior performance in both IF Eval and Pair-wise winning scores, with slightly lower accuracy on the Open LLM leaderboard. This result suggests that “Diverse Temp.” enhances the model’s constraint controllability while enhancing its general instruction-following capabilities compared to “Single Temp.” On the contrary, when fixing the rule templates to one single template, potential overfitting to the template might occur and thus negatively influence the model performances.
Effect of Rule Numbers: “Max Rule = $x$” represents the setting in which at most $x$ different rules can be sampled and utilized on each original data sample. In the augmentation process, a random value will be sampled in the range of $[0,x]$ as the number of rules in this sample. The examples in Figure 2 showcase the scenario when only one rule is implemented and examples in Figure 3 showcase the scenario when multiple rules are implemented. Compared to the baseline, nearly all settings show performance improvements across the three evaluation metrics. However, the IF Eval score initially increases, reaching its peak when “Max Rule = $3$”, before declining. The Pair-wise score, on the other hand, consistently decreases from “Max Rule = $1$” to “Max Rule = $5$”. These results suggest that applying too many rules to a single sample may impair the LLM’s capability, even though sampling and applying multiple rules can be beneficial when done in moderation, which might be because the original instruction becomes so complex.
Effect of Augmentation Rate: “Aug Rate = $x$” represents there is a probability of $x$ to apply our augmentation to each sample. It is observed that as the augmentation rate increases, performance improves on the IF Eval and mostly improves on the Pair-wise evaluation. This phenomenon indicates that increasing the augmentation rate primarily enhances the LLM’s constraint controllability, and it also has a positive impact on its general instruction-following capability. However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
Detailed Sub-Category Analysis: The detailed sub-category analysis can be found in Appendix C.
4 Conclusion
In this work, we proposed Rule-based Data Recycling (RuleR), which modified the original instructions and responses from an existing dataset by rule-defined constraints. The “recycled” data aims to enhance the LLMs’ capability to generate outputs fulfilling the constraints specified in the input, thereby improving the controllability of LLMs. RuleR took the first step of exploring rule-based data recycling, which can serve as a plug-and-play and easy-to-use method that converts any existing SFT datasets to new datasets for better controllability.
Limitations
Appendix A Experimental Setup
A.1 Implementation Details
The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. For training on Llama2-7B and Llama2-13B Touvron et al. (2023), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs. For training on Mistral-7B Jiang et al. (2023), the maximum learning rate is set to $1\times 10^{-5}$ with a warmup rate of $0.1$ for $2$ epochs. When utilizing our method, we run the augmentation process $3$/$2$ times to simulate the epochs of training. These augmented data are then mixed together and used for training $1$ epoch. All other configurations are kept the same as the baselines.
A.2 Training Datasets
We utilize 5 SFT datasets to evaluate the effectiveness of our method:
A.3 Evaluation Metrics
Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts. This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al. (2024); Li et al. (2023b); Sottana et al. (2023). We utilize test instruction sets from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023b), comprising 218 and 80 diverse human-curated instructions, respectively. Following the framework by Li et al. (2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy. To mitigate positional bias Ko et al. (2020); Wang et al. (2023a), responses are presented in two distinct sequences. Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction. One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions. The specific evaluation prompt is presented in the Appendix D.
Appendix B Examples
Examples of multi-rule augmentation are shown in Figure 3.
[Instruction]
Ensure the 2nd paragraph is enclosed in "-| |-" in the response.
[Response]
[Instruction]
Keep the word "way" bounded by "@ @" in the response if it appears.
Describe an example of a time you used to influence in a positive way
[Response]
Appendix C Detailed Sub-Category Analysis